4 times Harry's age equals 110 minus 3 times Grace's age minus 3 times Dave's age, 2 times Grace's age equals 76 minus 2 times Dave's age minus 3 times Harry's age, and 2 times Grace's age equals 68 minus 2 times Dave's age minus 2 times Harry's age. What is each person's age?

Answer by ankor@dixie-net.com(22740)

Write an equation for each statement
;
4 times Harry's age equals 110 minus 3 times Grace's age minus 3 times Dave's age,
4h = 110 - 3g - 3d
2 times Grace's age equals 76 minus 2 times Dave's age minus 3 times Harry's age,
2g = 76 - 2d - 3h
and 2 times Grace's age equals 68 minus 2 times Dave's age minus 2 times Harry's age.
2g = 68 - 2d - 2h
:
2g = 2g, therefore we can write the equation
76 - 2d - 3h = 68 - 2d - 2h
add 2d to both sides
76 - 3h = 68 - 2h
76 - 68 = -2h + 3h
8 = h
Harry is 8 yrs old
that is about all I can extract from this problem
What is each person's age?